OpenStudy (anonymous):

Write an equation of the line parallel to y=8x+1 that contains (-6,2)

OpenStudy (ccswims):

We know that when it says parallel it is saying keep the slope, so right away we know our slope is going to be positive 8

OpenStudy (ccswims):

We are first going to have to do point-slope form in order to write an equation, do you know how or do you need help with point-slope form?

OpenStudy (anonymous):

point slope is y-y1=m(x-x1) right?

OpenStudy (ccswims):

yes, so plugging in everything what will you get?

OpenStudy (anonymous):

y-2=8(x-1)?

OpenStudy (ccswims):

it wouldn't be -1

OpenStudy (ccswims):

it would be x+6, do you know how I got that?

OpenStudy (anonymous):

yeah i was using the wrong number in the equation

OpenStudy (ccswims):

ok :) so now you know its y-2 = 8(x+6) Now you have to distribute the 8 to x and 2

OpenStudy (anonymous):

actually the equation was all i had to do but thank you so much for the help

OpenStudy (ccswims):

ok cool np

OpenStudy (anonymous):

ok so this one i have to write an equation for the line that passes through (4.2) and (3,-2)

OpenStudy (ccswims):

do you mean (4,2)?

OpenStudy (anonymous):

yes

OpenStudy (ccswims):

so first you need to find the slope y2 - y1 over x2-x1

OpenStudy (anonymous):

so it would be m=-2-2 over 3-4 right?

OpenStudy (ccswims):

yes

OpenStudy (ccswims):

just simplfy

OpenStudy (ccswims):

like what is -2-2 and what is 3-4

OpenStudy (ccswims):

as in sum

OpenStudy (anonymous):

so -4over 1

OpenStudy (ccswims):

it would be -4/-1 since both numbers are negative they cancel each other out, so the real slope is just 4 (4/1 simplified is 4)

OpenStudy (anonymous):

okay that makes since

OpenStudy (ccswims):

Now do point-slope form again

OpenStudy (anonymous):

well all i had to find was the equation

OpenStudy (ccswims):

ok nvm the equation should be y=mx+b "b" being y-intercept, to find the y-intercept, which is 14

OpenStudy (ccswims):

so y=4x + 14

OpenStudy (ccswims):

or sorry y = 4x - 14

OpenStudy (ccswims):

^^that is your answer